Practice question: need help understanding how the answer to (part a is 15) and (part b is 150)

A: Suppose we have an economy for which G = 300, T = 240, S = 80, I = 45, and imports = 40. Exports must be?

B: If S = 200, T = 700, G = 950, and NX = -200, this makes domestic investment?
Answer #1

A) use the rule

saving - investment = net exports.

In this case private saving is 80. Public saving is T - G = -60. This gives the National savings at 80+(-60) = 20. Now we have saving - investment= exports - imports

20 - 45 = X - 40

X = 60-45 = 15

B) in this case public saving is - 250. Private saving is 200 which means National savings is -250+200 = -50. Domestic investment = savings - net exports or -50 - (-200) = 150